OpenStudy (anonymous):

Use two variables and two equations. Solve by any method you choose. Write your answer in the appropriate form. To earn money for college, Susan is making and selling earrings. Her weekly costs for advertising and phone calls are $36 and each pair of earrings cost $1.50 to produce. If Susan sells the earrings for $6 per pair, how many pairs must she sell per week to break even?

OpenStudy (mathmate):

You do not require a system of equations to solve this. Use arithmetic. She makes 6-1.5=4.5 per pair of earrings. To cover fixed costs of $36, she needs to sell (36/4.5)=8 pairs. You can always solve by the complicated way if you want/have to.

OpenStudy (anonymous):

???? so i don't have to solve just write it out

OpenStudy (mathmate):

I think you need to use two variables, set up two equations, and solve it by any method you prefer. Do you need help doing any of the above?

OpenStudy (anonymous):

yes plz

OpenStudy (mathmate):

Let x=number of pairs of earrings she sells per week y=profit per pair then y=6-1.50 x*y-36=0 I'll leave it to you to solve the equations.